OpenStudy (anonymous):

Solve this question without a calculator: log4 20 − log4 45 + log4 144

OpenStudy (anonymous):

answer will be 3 i think do like this first apply loga - logb=log(a/b) to log4 20 and log4 45 then apply loga + logb=log(ab) to the remaining two terms you should get your answer

OpenStudy (yttrium):

It can be rewritten like this \[\log _{4}[\frac{ 20*144 }{ 45 }]\] And that will result to \[\log _{4}64\] which is equal to 3